# Exploring the Nostr Ecosystem: A Study of Decentralization and Resilience

### v1

@@ Abstract.
Nostr is an open decentralized social network launched in 2022. From a user’s perspective, it is similar to a micro-blogging service like Twitter. However, the underlying infrastructure is very different, and Nostr boasts a range of unique features that set it apart. Nostr introduces the concept of relays, which act as open storage servers that receive, store, and distribute user posts. Each user is uniquely identified by a public key, ensuring authenticity of posts through digital signatures. Consequently, users are able to securely send and receive posts through various relays, which frees them from single-server reliance and enhances post availability (e.g., making it more censorship resistant). The Nostr ecosystem has garnered significant attention, boasting 4 million users and 60 million posts in just 2 years. To understand its characteristics and challenges, we conduct the first large-scale measurement of the Nostr ecosystem, spanning from July 1, 2023, to December 31, 2023. Our study focuses on two key aspects: Nostr relays and post replication strategies. We find that Nostr achieves superior decentralization compared to traditional Fediverse applications. However, relay availability remains a challenge, where financial sustainability (particularly for free-to-use relays) emerges as a contributing factor. We also find that the replication of posts across relays enhances post availability but introduces significant overhead. To address this, we propose two design innovations. One to control the number of post replications, and another to reduce the overhead during post retrieval. Via data-driven evaluations, we demonstrate their effectiveness without negatively impacting the system.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Social media platforms such as Facebook and Twitter rank among the most frequented websites globally. However, despite their success, they face increasing criticism (Raman et al., 2019). For example, the dominant, uncompetitive nature of these platforms has granted them unrestricted discretion over permissible activities, triggering regulatory concerns and sparking widespread debates (Satariano, 2021; Frier et al., 2021).
@@
In response, there has been growing movement of activists, technologists and researchers working on the ”Decentralized Web” (DW). This encompasses technologies that broadly aim to foster increased transparency, openness, and democracy on the web (Nystedt, 2018). In particular, the ”Fediverse”, composed of server-based federated services such as Mastodon, Pleroma, and other Twitter-like micro-blogging services, rank among the most widely used applications within the DW (Zia et al., 2022). These platforms deconstruct their service offerings into independent servers (instances) that anyone can easily create. In the simplest scenario, these instances permit users to register and interact locally. However, they also enable cross-instance interaction (i.e., federation). This distributed structure offers numerous benefits. For instance, it provides more transparent data ownership, makes censorship more difficult. Further, the lack of centralisation could potentially enhance the overall system’s resilience against technical, legal, or regulatory attacks.
@@
Despite the benefits, these platforms also suffer from key limitations (Raman et al., 2019; Hassan et al., 2021). User identities are tied to a server controlled by an administrator. Thus, instance owners can ban users (just as Twitter can), and they also have the power to block other instances, thereby imposing censorship. Perhaps more problematic is that there are no clear incentives for running these servers. Consequently, they are often managed by enthusiasts who may operate them with more passion than professional polish. For example, a recent study found that 11% of instances are inaccessible for more than half of the time (Raman et al., 2019). This situation puts users at the mercy of a single individual administrator’s whims, which can sometimes be even more oppressive than a large corporation like Twitter (Rozenshtein, 2023).
@@
To address the limitations of the Fediverse, a fundamentally different decentralized web platform was launched in 2022: Nostr (Notes and Other Stuff Transmitted by Relays) (Nostr, 2023a). Nostr is a decentralized system that is built on a series of open protocols capable of establishing a global social network in a way that detaches users from any specific server (which could later become unavailable or censor the users). From a user’s perspective, it exhibits similar functions to a micro-blogging service like Twitter, although the actual implementation differs significantly. The core innovation is the introduction of relays. These operate as open storage servers, loosely akin to the server instances in the Fediverse. Their only responsibility is to receive posts from users, store them, and distribute them to other users. However, unlike in the Fediverse a relay does not communicate with other relays; it interacts solely with users. Further, users can push posts to any (including multiple) relays, thereby decoupling the fixed association between a user and a server. Thus, even if individual relays fail (or censor content), the posts remain live. Others wishing to retrieve posts can simply connect to any (including multiple) of these relays to collect the posts. To ensure integrity, each user is uniquely identified by a public key, and each message they create is digitally signed to ensure authenticity. These unique features have garnered significant attention from users, and fostered numerous applications built atop of Nostr. To date, the Nostr ecosystem encompasses more than 600 active relays (Nostr.Watch, 2023), more than 4 million users, and more than 60 million posts (Nostr.Band, 2023).
@@
As the Nostr ecosystem continues to grow and gain traction, we argue that it is crucial to better understand the characteristics and challenges of the decentralized Nostr ecosystem. We are particularly curious about whether Nostr encounters similar challenges of the Fediverse, such as poor relay (instance) availability, post availability, and the natural tendency of centralization (Raman et al., 2019). We are also curious about any novel challenges that may uniquely emerge in Nostr. We conjecture that the replication of posts across relays significantly increases overheads for storage and relay-user communications. To date, these challenges are not understood and no explicit and deterministic solutions exist.
@@
With this in mind, we introduce the first large-scale measurement of the Nostr ecosystem. We assemble a dataset of 17.8 million posts, 1.5 million unique pubkeys (users), and 712 relays, spanning the time period from 2023-07-01 to 2023-12-31. Using this dataset, we focus on two key aspects.
@@
First, we study the Nostr relays, regarding the availability and the degree of decentralization. Through the relay-based paradigm, we show that Nostr has accomplished significantly better decentralization compared to prior Fediverse applications (§4.1 §4.2). The users and posts are highly distributed across the relays. Most (93%) posts can be found in multiple relays, with 178 relays (25%) respectively host a significant proportion (more than 5%) of the posts. The relays are distributed over 44 countries and 151 autonomous systems. Thus, surprisingly, no single country or autonomous system hosts more than 25% of the relays, and the majority have only 1 or 2 relays each. However, we find that the unstable availability of relays is still a challenge, akin to prior Fediverse applications (Raman et al., 2019)4.3). We show that 20% of the relays experience downtime for more than 40% of the time. Our further investigation suggests that financial sustainability may be one reason for this, especially for free-to-use relays (§4.4). Although Nostr features build-in support for Bitcoin, enabling users to tip or donate to relay operators, our estimation shows that 95% of the free-to-use relays cannot cover their operational cost from this.
@@
We next focus on the replication and availability of posts. A unique innovation of Nostr is the ability to place posts across multiple relays. This, however, impose trade-off between improved post availability (as well as resilience to censorship) and lower replication overhead. We show that post availability on Nostr significantly surpasses that of Fediverse applications: even when the top 30 relays with the highest number of users or posts are down, over 90% of the posts remain accessible in Nostr (§5.2). Initially, this suggests that Nostr has accomplished its core anti-censorship and resilience goal. However, the overhead of post replications is large. We find 616M post replications for the 17.8M posts, which means, on average, a post is replicated across 34.6 relays (§5.1 §5.3). Another challenge stemming from post replications is that due to the Nostr protocol, users may retrieve the same post multiple times from different relays (§5.4). Indeed, Our evaluation indicates that 98.2% of these retrievals are redundant, leading to a wasteful consumption of bandwidth.
@@
These results motivate us to study if it is possible to reduce the overhead of post replications. Consequently, we propose a new feature in post publication for Nostr clients (§6.1 §6.2). Our evaluation demonstrates that this feature can effectively reduce 380 million post replications while preserving the same level of post availability. Moreover, we propose another feature in post retrieval (§6.3 §6.4). According to our evaluation, this feature can effectively decrease the wasteful retrievals from 98.2% to 11.7%, while ensuring that overall 99.4% of post can be found and retrieved.
The paper makes three key contributions:
@@
    We compile the first large-scale dataset of the Nostr ecosystem, comprising 17.8 million posts, 616 million post replications, 1.5 million users, and 712 relays.
@@
    Through our comprehensive dataset, we observe significant differences between Nostr and the Fediverse, as decentralized social networks. While we highlight that Nostr is more decentralized with improved post availability, it comes at the cost of notable overhead.
@@
    In response, we propose two design innovations. The first aims to control the number of post replications, while the second aims to minimize the overhead during post retrieval. We demonstrate that they are effective without causing negative impact on the system.
@@ 2. Nostr Primer
Nostr Overview.  Nostr is a decentralized social networking system built on a series of open protocols. From a user’s perspective, it can be seen as a micro-blogging service similar to Twitter or Mastodon. However, the underlying architecture is very different. Nostr is composed of two key elements: relays and clients. A relay operates as a server, receiving messages from certain users and forwarding them to others (upon request) in an asynchronous manner. Users interact with relays, sending and receiving messages through a client. This client can be any application (mobile, desktop, web, etc.) that implements the Nostr client protocol.
Any individual can operate a relay, akin to the Fediverse where anyone can set up an instance. Similarly, end users can select any client implementation, or even build their own compatible client. For context, Figure 1 offers the screenshots of two Nostr clients. [^1] Screenshot of Iris is retrieved from https://www.alza.cz/nostr-decentralizovana-sit-pro-svobodnejsi-sdileni-obsahu-na-internetu, and screenshot of Damus is retrieved from https://github.com/damus-io/damus. When using these clients, a user can create a post, which is then pushed to the relays of their choice. Similarly, the clients can connect to relays to download posts. Further details about this process are provided below.
User Account.  To begin, the user needs an account, which is a public key. Nostr utilizes asymmetric encryption, identifying each user by their respective public key. The user generates a key pair and provides it to the client to facilitate interactions with relays. Crucially, as long as the user maintains their private key, they have ultimate ownership of the account, unlike platforms like Twitter or Mastodon where the account can be manipulated or deleted by system admins. This flexibility also enables the user to seamlessly switch between relays while maintaining the same account. This makes it notably different to prior platforms, which retain ultimate account ownership.
Event.  For data transmission, the Nostr protocol specifies the event concept, represented in a distinct JSON schema. Nostr defines various kinds of events for different purposes, such as user profiles (kind-0), posts (kind-1), following lists (kind-3). A post (kind-1) is the fundamental and most commonly used event, functioning similarly to a tweet on Twitter or a toot on Mastodon. More details can be found in the protocol specification (Nostr, 2023b).
@@
Connecting to Relays.  WebSocket (WebSocket, 2023) is used as the underlying protocol to connect to relays. The Nostr protocol dictates that clients should initiate a single websocket connection to each relay and employ it for all forms of communication with this relay.
@@
Receiving from Relays.  To receive posts from a relay, the client must create a subscription and send it to the relays. Each subscription includes a filter that specifies the types of events the client wishes to receive. The relay then forwards the desired events to the client. For instance, if user $F$ follows user $A$, $F$’s client will create a subscription with a filter for $A$’s recent events. Thus, client $F$ will then connect to the relay and upload the filter. The relays will then forward $A$’s events to the client and continue to forward new events from $A$ once they are received.
Zaps.  Zaps function as monetary tips that one user can send to another, which is also used to fund relays. A user can opt to link their account with a Bitcoin wallet address, thereby enabling them to send or receive Bitcoins in the form of zaps. A successful zap transaction results in the publication of a zap receipt within the relays as a kind-9735 event. More details can be found in (Nostr, 2023d).
@@ 3. Dataset
Recall, we focus on two key aspects of the Nostr ecosystem: relays and post replications across relays. To measure these, we gather two datasets. The first is the Nostr relay dataset, which includes general information, deployment, and availability data of the relays. The second is the event dataset, which contains the events in the relays. This dataset helps us gauge the scale and popularity of the relays, as well as measure post replications and availability.
@@ 3.1. Relay Dataset
Relay Selection.  We begin by compiling a list of relay URLs and IP addresses from the nostr.watch website, which contains an index of relays around the world. We extract the relays from this site on a daily basis from 2023-10-01 to 2023-12-31. We collect the list of 911 relays from nostr.watch that have been online at least once during this period. We successfully retrieve event data from 712 of them. We cannot obtain event data from the remaining 199 relays due to three reasons: (i) some relays (58%) have no events published, indicating that they are either not actively used or solely set up for testing purposes; (ii) certain relays (41%) remain online for only a short period before going constantly offline; and (iii) some relays (1%) require payment for accessing their data. Consequently, all analyses conducted in this study exclude these relays from which we cannot obtain event data.
Relay Information.  Relays provide server metadata (a relay information document) to clients. This document provides clients with information on server capabilities, administrative contacts, and various server attributes, such as supported NIPs, content limitations, community preference, and payment information. We provide a sample of relay information document in Appendix A.1, and more detail can be found in the protocol specification (Nostr, 2023c).
This information also allows us to separate the relays into free relays and paid relays. In the case of a paid relay, users are required to pay a certain amount of bitcoin to use it. While most paid relays only charge for posting to the relay, a small number charge for reading (we exclude them because we cannot crawl them). Of the 712 relays in our dataset, 625 are free and 87 are paid.
Autonomous Systems.  We then query IP-API with the IP address of the relay to get the Autonomous Systems (AS) hosting the relay, as well as the geographical location (country and city) of the relay server.
Relay Availability.  We utilize the API provided by nostr.watch to monitor the availability of relays. Every 15 minutes, we request information from the API and receive a list of online relays. Our monitoring period is from 2023-10-01 to 2024-12-31. This allows us to map the exact downtime of each relay.

### v2

# An Empirical Analysis of the Nostr Social Network: Decentralization, Availability, and Replication Overhead

@@ Abstract.
Nostr is a decentralized social network launched in 2022, emphasizing high availability and censorship resistance. Since launching, it has gained substantial attention, boasting over 100 million posts. From a user’s perspective, it is similar to a micro-blogging service like Twitter. However, the underlying systems infrastructure is very different, and Nostr boasts a range of unique features that set it apart. Nostr introduces the concept of relays, which act as open storage servers that receive, store, and distribute user posts. Each user is uniquely identified by a public key, ensuring authenticity of posts through digital signatures. Users are able to securely replicate and retrieve posts through multiple relays, which frees them from single-server reliance and enhances post availability, thereby attempting to make Nostr censorship resistant. However, this aggressive design also presents challenges, such as the overhead required for extensive post replication and the difficulty in obtaining a global view of post replication locations, which remain unexplored or unaddressed. This necessitates a thorough understanding of the Nostr ecosystem; therefore, we conduct the first large-scale study on this topic. Our study focuses on two key aspects: Nostr relays and post replication strategies. We find that Nostr achieves superior decentralization compared to traditional Fediverse applications. However, relay availability remains a challenge, where financial sustainability (particularly for free-to-use relays) emerges as a contributing factor. We also find that the replication of posts across relays enhances censorship-resistance but introduces significant overhead. To address this, we propose two improvements: one to control the number of post replications, and another to reduce the overhead during post retrieval. Via a data-driven evaluation, we demonstrate their ability to reduce overhead without negatively impacting post availability under the simulated scenarios.
Nostr, Social Network, Decentralization, Measurement, Censorship
$c$s: Networks Network measurement
@@ 1. Introduction
Censorship resistance is a major research topic. Numerous approaches have been proposed (Khattak et al., 2016), including refraction networking (Bocovich and Goldberg, 2016), censorship-resistant CDNs (Zolfaghari and Houmansadr, 2016), packet manipulators (Bock et al., 2019), and WebRTC-based tools (Barradas et al., 2020; Figueira et al., 2022; Bocovich et al., 2024). Recently, censorship resistance in social media platforms has garnered particular attention (Huang, 2024; Shashlin, 2023). This has been driven by the monopolistic nature of traditional online social media platforms (such as Facebook and Twitter/X), which has given them significant control over permissible activities, raising growing concerns about censorship and freedom of speech (Satariano, 2021; Frier et al., 2021).
@@
Consequently, there has been a growing movement of technologists and researchers working on decentralized social networks that aim to foster increased transparency, openness, and democracy. The “Fediverse”, composed of server-based federated services, such as Mastodon, rank among the most widely used applications for decentralized social networks (Zia et al., 2022). These platforms deconstruct their service offerings into independent servers (instances) that anyone can easily setup. In the simplest scenario, these instances permit users to register and interact locally. However, they also enable cross-instance interaction (i.e., federation). This distributed structure offers numerous benefits. For instance, it provides more transparent data ownership, and potentially enhances the overall system’s resilience against technical or legal attack.
@@
Despite the benefits, these platforms suffer from key limitations (Raman et al., 2019; Hassan et al., 2021). User identities are tied to a single server controlled by an administrator. Thus, instance owners can ban users (just as Twitter/X can), delete content, and even block other servers, thereby imposing censorship. Perhaps more problematic is that there are no clear incentives for running these servers. Consequently, it can be challenging for part-time enthusiasts to resource such activities. For example, a recent study found that 11% of instances are inaccessible for more than half of their lifespan (Raman et al., 2019). This situation puts users at the mercy of a single administrator’s personal preferences, which sometimes have been more oppressive than large corporations like Twitter/X (Rozenshtein, 2023).
To address the limitations of the Fediverse, a fundamentally different decentralized web platform was launched in 2022, receiving significant financial backing from Jack Dorsey: Nostr (Notes and Other Stuff Transmitted by Relays) (Nostr, 2023). From a user’s perspective, Nostr exhibits similar functions to a micro-blogging service like Twitter/X, although the actual implementation differs substantially. Nostr is a decentralized system that is built on a series of open protocols capable of establishing a global social network in a way that detaches users from any specific server. Publicly, Nostr states that this approach is designed with one key requirement in-mind: attaining censorship resistance against administrators removing content or blocking specific users.
@@
The core innovation is the introduction of relays, which operate as open storage servers, loosely akin to the server instances in the Fediverse. Their only responsibility is to receive posts from users, store them, and distribute them to others upon request. However, unlike in the Fediverse, a relay does not communicate with other relays; it interacts solely with users. Further, users can push posts to any (including multiple) relays, thereby decoupling the fixed association between a user and an individual server. Thus, even if individual relays censors content (or simply fail), the posts remain live. Note, this may also offer greater censorship resistance against other forms of censorship (e.g., state-level firewalls), but Nostr does not explicitly strive to address such scenarios. To ensure integrity, each user is uniquely identified by a public key, and each message is digitally signed. These unique features have garnered significant attention from users, and fostered numerous applications built atop of Nostr. To date, the Nostr ecosystem encompasses more than 600 active relays (Nostr.Watch, 2023), more than 9 million users (with a profile and contact list), and 100 million posts (Nostr.Band, 2023).
@@
As the Nostr ecosystem continues to grow, we argue that it is crucial to better understand its characteristics and the associated challenges. We are particularly curious about whether Nostr encounters similar resilience challenges to that of the Fediverse, such as poor relay (instance) availability, post availability, and the natural tendency towards centralization (Raman et al., 2019). Critically, we wish to test how effectively Nostr achieves the goal of censorship resistance, and if this comes at the cost of high overhead. For example, the replication of posts may significantly increase storage overhead and relay-to-user communication costs. To date, these challenges are not understood and we lack any deterministic solutions.
@@
With this in mind, we present the first large-scale study of the Nostr ecosystem. We assemble a dataset of 17.8 million posts, 1.5 million unique pubkeys (users), and 712 relays, spanning the time period from 2023-07-01 to 2023-12-31. Using this dataset, we focus on two key aspects.
@@
First, we study Nostr from the perspective of relays, inspecting their availability, reliability, and the degree of decentralization. Through the relay-based paradigm, we show that Nostr has accomplished significantly greater decentralization compared to prior Fediverse applications (§4.1, §4.2). The users and posts are highly distributed across the relays, attaining significant censorship resilience. Most (93%) posts can be found in multiple relays, with 178 relays (25%) individually hosting a significant proportion (more than 5%) of global posts. The relays are distributed over 44 countries and 151 autonomous systems (ASes). Thus, no single country or autonomous system hosts more than 25% of the relays, and the majority have only 1 or 2 relays each, making censorship attacks far more difficult. However, we find that the unstable availability of relays is still a challenge (§4.3). We show that 20% of the relays experience downtime for more than 40% of the measurement period. Further investigation suggests that financial sustainability may be a key reason for this, especially for free-to-use relays (§4.4). Although Nostr offers support for user donation, our estimates shows 95% of the free-to-use relays cannot cover their operational cost from this alone.
@@
We next focus on the replication and availability of posts. A unique innovation of Nostr is the ability to push posts across multiple relays. This, however, imposes a trade-off between improved post availability (i.e., censorship resistance) and lower replication overhead. We show that post availability on Nostr significantly surpasses that of Fediverse applications: even when the top 30 relays with the highest number of users or posts are down, over 90% of the posts remain accessible (§5.2). Initially, this suggests that Nostr has accomplished its core anti-censorship and resilience goal. However, the overhead of post replications is large. We find 616M post replications for 17.8M posts. This means, on average, a post is replicated across 34.6 relays (§5.1, §5.3). Another challenge stemming from post replications is that, due to the Nostr protocol design, in many cases, users must retrieve the same post multiple times from different relays. Indeed, our evaluation indicates that 98.2% of these retrievals are redundant, leading to extremely wasteful consumption of bandwidth  (§5.4).
@@
These results motivate us to study if it is possible to reduce the overhead of post replications. Consequently, we propose a new feature in post publications for Nostr clients (§6.1) that automatically limits the number of replications and strategically selects relays for these replications. Our evaluation demonstrates that this feature can effectively reduce 380 million (61.7%) post replications while preserving the same level of post availability and censorship resistance under the simulated scenarios of relay-level and AS-level failures. Moreover, we propose another feature in post retrieval for Nostr clients (§6.2) to minimize wasted bandwidth by optimistically querying only a subset of known relays. According to our evaluation, this feature effectively decreases wasteful retrievals by 93%, while ensuring that 99.4% of the posts can be retrieved. This paper makes three key contributions:
@@
    We perform the first large-scale systems study of the Nostr ecosystem, covering 17.8 million posts, 616 million post replications, 1.5 million users, and 712 relays.
@@
    We confirm that Nostr achieves high content availability, making censorship extremely difficult. However, while we highlight that Nostr is more decentralized with improved post availability and censorship resistance, we show that it comes with significant overhead.
@@
    In response, we propose two design improvements: One aims to control the number of post replicas, and one aims to minimize the overhead during post retrieval. We show that in our data-driven simulation scenarios, they are effective without causing negative impact on the censorship resistance.
@@ 2. Nostr Primer
Nostr Overview.  Nostr is a decentralized social networking system built on a series of open protocols. From a user’s perspective, it can be seen as a micro-blogging service similar to Twitter/X or Mastodon. However, the underlying architecture is very different. Nostr is composed of two key elements: relays and clients. A relay operates as a server, receiving messages from certain users and forwarding them to others (upon request) in an asynchronous manner. Users interact with relays, sending and receiving messages through a client. This client can be any application (mobile, desktop, web) that implements the Nostr client protocol.
Any individual can operate a relay, akin to the Fediverse where anyone can set up an instance. Similarly, end users can select any client implementation, or even build their own compatible client. For context, Figure 11 in the Appendix offers the screenshots of two Nostr clients. When using these clients, a user can create a post, which is then pushed to the relays of their choice. Similarly, the clients can connect to relays to download posts. Further details about this process are provided in the following paragraphs. The design aims to achieve strong censorship resistance by distributing post replications across multiple relays. This makes it difficult for attacks or failures to affect all relays simultaneously. As long as at least one relay continues to host the post, it remains available, thereby ensuring censorship resistance.
User Account.  To begin, the user needs an account, which is a public key independent from any relay. Nostr utilizes asymmetric encryption, identifying each user by their respective public key. The user generates a key pair and provides it to the client to facilitate interactions with relays. Crucially, as long as the user maintains their private key, they have ultimate ownership of the account, unlike platforms like Twitter/X or Mastodon where the account can be manipulated or deleted by system admins. This flexibility also enables the user to seamlessly switch between relays while maintaining the same account. This makes it notably different to prior platforms, which retain ultimate account ownership.
Event.  The event is defined as a specific JSON schema, which serves as the fundamental data structure for data transmission in Nostr. Nostr then defines various kinds of events for different purposes, such as user profiles (kind-0), posts (kind-1), following lists i.e., lists of other users followed by the user (kind-3). A post (kind-1) is the core and most common event, functioning similarly to a tweet on Twitter/X. Together, these events facilitate the operation of a social network. More details can be found in the protocol specification (Nostr-NIP-01, 2023).
@@
Connecting to Relays.  Nostr uses WebSockets (WebSocket, 2023) to connect to relays. The Nostr protocol dictates that clients should initiate a single WebSocket connection to each relay and employ it for all forms of communication with this relay.
@@
Receiving from Relays.  To receive posts from a relay, the client must create a subscription and send it to the relays. Each subscription includes a filter that specifies the kinds of events the client wishes to receive. The relay then forwards the desired events to the client. For instance, if user $F$ follows user $A$, $F$’s client will create a subscription with a filter for $A$’s recent events. Thus, client $F$ will then connect to the relay and upload the filter. The relays will then forward $A$’s events to the client and continue to forward new events from $A$ once they are received.
Zaps.  Zaps function as monetary tips that one user can send to another, which is also used to fund relays. A user can opt to link their account with a Bitcoin wallet address, thereby enabling them to send or receive Bitcoins in the form of zaps. A successful zap transaction results in the publication of a zap receipt within the relays as a kind-9735 event. More details can be found in (Nostr-NIP-57, 2023).
@@ 3. Dataset
We focus on two key aspects of the Nostr ecosystem: relays and post replications across relays. To measure these, we gather two datasets. The first is the Nostr relay dataset, which includes general information, deployment, and availability data of the relays. The second is the event dataset, which contains the events in the relays. This dataset helps us gauge the scale and popularity of the relays, as well as measure post replications and availability. See §A.1 for an ethics discussion.
@@ 3.1. Relay Dataset
Relay Selection.  We begin by compiling a list of relay URLs and IP addresses from the nostr.watch website which is widely utilized within the Nostr ecosystem. It maintains a comprehensive index of relays around the world by implementing Nostr’s relay discovery and monitoring protocol (Nostr-NIP-66, 2023). We extract the relays from this site on a daily basis from 2023-07-01 to 2023-12-31. We collect the list of 911 relays that have been online at least once during this period. We successfully retrieve event data from 712 of them. We cannot obtain event data from the remaining 199 relays due to three reasons: (i) some relays (58%) have no events published, indicating that they are either not actively used or solely set up for testing purposes; (ii) certain relays (41%) remain online for only a short period before going constantly offline; and (iii) some relays (1%) require payment for accessing their data. Consequently, all analyses conducted in this study exclude these relays from which we cannot obtain event data.
Relay Information.  Relays provide server metadata (a relay information document) to clients. This document provides clients with information on server capabilities, administrative contacts, and various server attributes, such as content limitations, community preference, and payment information. We provide an example relay information document in Appendix A.3, and more details can be found in the protocol specification (Nostr-NIP-11, 2023). This information also allows us to separate the relays into free relays and paid relays. In the case of a paid relay, users are required to pay a certain amount of bitcoin. While most paid relays only charge for posting to the relay, a small number charge for reading (we exclude them because we cannot crawl them). Of the 712 relays in our dataset, 625 are free and 87 are paid.
Autonomous Systems.  We then query IP-API (IP-API, 2023) with the IP address of the relay to get the AS hosting the relay, as well as the geographical location (country) of the relay server.
Relay Availability.  To monitor the availability of relays, we utilize the feature provided by nostr.watch. The nostr.watch website conducts frequent periodic measurement to test the availability of the relays. Every 15 minutes, we take a snapshot of the availability status of relays reported by nostr.watch. Our monitoring period is from 2023-10-01 to 2023-12-31. This allows us to map the exact downtime of each relay.